OpenStudy (anonymous):

What is the next number in the pattern: 1, -2, 5, 2, 9, 6, 13, …?

OpenStudy (anonymous):

10

OpenStudy (tkhunny):

-2 34 \(\sqrt{2}\) Take your pick. Any of them can be next.

OpenStudy (anonymous):

10

OpenStudy (tkhunny):

18 \(-\pi\) 3.78 Take your pick. Any of them can be the next value.

OpenStudy (anonymous):

The pattern is +4.

OpenStudy (anonymous):

the pattern is not +4

OpenStudy (anonymous):

Hmmm... Im learning this in school right now... but i may be wrong.

OpenStudy (tkhunny):

"A" pattern is "-3 followed by +7". A common way to proceed is call a "Difference". Just subtract each successive pair. 1, -2, 5, 2, 9, 6, 13, -2 - 1 = -3 5 - (-2) = +7 2 - 5 = -3 9 - 2 = +7 6 - 9 = -3 13-6 = +7 It is reasonable to assume the next will be another -3, giving 10. But, this is only one of infinitely many possible solutions to this question.

OpenStudy (anonymous):

Oh, because this is what I was thinking... 

OpenStudy (tkhunny):

That's fine. Every other one is a +4.

OpenStudy (anonymous):

i understand later

OpenStudy (anonymous):

36

OpenStudy (anonymous):

Thank you for all the help guys :)

OpenStudy (anonymous):

No problem